### v5

# Reviewing Uses of Regulatory Compliance Monitoring

Abstract
Organizations need to manage numerous business processes for delivering their services and products to customers. One important consideration thereby lies in the adherence to regulations such as laws, guidelines, or industry standards. In order to monitor adherence of their business processes to regulations – in other words, their regulatory compliance – organizations make use of various techniques that draw on process execution data of IT systems that support these processes. Previous research has investigated conformance checking, an operation of process mining, for the domains in which it is applied, its operationalization of regulations, the techniques being used, and the presentation of results produced. However, other techniques for regulatory compliance monitoring, which we summarize as compliance checking techniques, have not yet been investigated regarding these aspects in a structural manner. To this end, this work presents a systematic literature review on uses of regulatory compliance monitoring of business processes, thereby offering insights into the various techniques being used, their application and the results they generate. We highlight commonalities and differences between the approaches and find that various steps are performed manually; we also provide further impulses for research on compliance monitoring and its use in practice.
keywords:
Conformance Checking, Compliance Monitoring, Regulations, Literature Review
Acknowledgements We thank Ilona Bogatinovska and Tom Knoche for their contributions towards this article.
Funding Information Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation), grant no. 465904964.
1 Introduction
In order to deliver their services and products to customers, organizations must manage a broad range of business processes [dumas2013fundamentals]. These business processes can be understood as structured and repeatable sets of activities designed to achieve specific business objectives [Weske19]. In practice, these business processes are usually subject to regulations, such as laws, industry standards, or guidelines [governatoriJourneyBusinessProcess2009]. These regulations originate from the processes environment and govern their execution by imposing constraints upon them. For businesses to achieve organizational success, it is important to comply with relevant regulations [lopezBusinessProcessCompliance2020], as not doing so could bring about unintended consequences, such as fines or loss of reputation [klessascheckReviewingConformance2024].
When aiming to ensure that all relevant regulations are followed, organizations usually face two options: They can either assess the compliance of their business processes when they are designed [hashmiAreWeDone2018] (so that an implementation is automatically compliant), or monitor their compliance during and after execution [groefsemaUseConformanceCompliance2022]. Notably, compliance monitoring brings the possibility of designing process execution in a more flexible manner, while still ensuring compliance, in comparison to a process that is compliant by design but must be strictly adhered to during execution to not violate compliance. One particular technique, conformance checking [carmona2018conformance], has been developed in the fields of business process management (BPM) and process mining, which has been identified as a useful technique for regulatory compliance monitoring [klessascheckReviewingConformance2024]: With conformance checking, the conformance of real-world process execution to desired process behaviour can be assessed in a data-driven manner.
Existing research has already investigated conformance checking uses for regulatory compliance monitoring [klessascheckReviewingConformance2024]; however, other techniques for checking compliance, such as those based on complex event processing also exist and can be applied for monitoring business process compliance [lyComplianceMonitoringBusiness2015]. While a wide range of research on applications of compliance monitoring has been conducted (e.g., case studies that use conformance checking [lemosUsingProcessMining2011], contributions that investigate the potential for regulatory compliance monitoring with conformance checking more abstractly [klessascheckUnlocking2024]), research currently does not provide a general overview of usages of various techniques available for monitoring regulatory compliance of business processes. To provide such an overview, thereby increasing awareness of existing approaches and identifying the potential for further contributions, we conduct a systematic literature review. In doing so, this work aims to answer the following research questions:
  1. RQ1:
    What are existing peer-reviewed studies that practically apply compliance monitoring techniques to assess the regulatory compliance of business processes?
  2. RQ2:
    What characterizes the actual compliance monitoring approaches in the contributions?
  3. RQ3:
    What are, based on the characterization of existing studies, potential areas of further investigation for the application of compliance monitoring techniques?
This paper is based on a previous study, where we reviewed conformance checking uses for regulatory compliance monitoring purposes [klessascheckReviewingConformance2024]. Extending our focus, we now additionally investigate usages of further techniques that exist in the research space of regulatory compliance and BPM in our review, and take further aspects such as the origin and availability of process execution data used in the investigated studies into account. We also describe our research method in more detail and, in light of our broadened scope, extend the results and discussion provided herein.
The remainder of this article is organized as follows: In Sec. 2, we provide background on compliance monitoring in general and conformance checking in particular, as well as related work. In Sec. 3 we present our research method for identifying and synthesizing relevant studies. In Sec. 4, we provide our analysis of the relevant studies, while we discuss our findings and the overall contributions in Sec. LABEL:sec:discussion. We conclude in Sec. LABEL:sec:concl, where we also present potential for future work.
2 Background and Related Work
In the following, we provide background on compliance monitoring, and discuss related work that has investigated these techniques and their applications.
2.1 Compliance Monitoring
As discussed above, it is necessary for businesses to assess whether their behaviour complies with the regulations to which they are subject. This act of ensuring that business processes do not violate the regulations relevant for them is called regulatory compliance checking [groefsemaUseConformanceCompliance2022]. One existing type of approach, being regulatory compliance monitoring (also known in the literature as run-time regulatory compliance checking [groefsemaUseConformanceCompliance2022]), aims at monitoring process executions to assess their compliance with regulations [vanbeestCrossInstanceRegulatoryCompliance2023]. This type of approach contrasts with traditional design-time regulatory compliance checking [groefsemaUseConformanceCompliance2022], in which the process model itself is analysed for violations of and compliance with regulations. However, processes can in practice be executed in ways that differ from the process model, thus giving relevancy to compliance monitoring during run-time [hashmiAreWeDone2018]. Note that while the compliance of business processes with business rules, that is, rules which embody organizational practice or policy [dumas2013fundamentals, Ch.4], can be assessed with compliance monitoring, we here focus specifically on regulatory compliance. This means assessing whether business processes comply with what law or rules ask or order them to do [groefsemaUseConformanceCompliance2022]. While other (system) requirements exist that need to be checked against business process models and executions (such as user requirements and design properties [groefsemaUseConformanceCompliance2022]), the importance of regulations that impose constraints on the behaviour of business processes, and the need of proving that these constraints are adhered to, motivates the focus of this study. Here, we concentrate on regulatory compliance of business processes specifically, instead of verifying the behaviour of a system and its implementation against a business process model more generally.
In research, various techniques for regulatory compliance monitoring have been identified as useful, among them conformance checking [klessascheckReviewingConformance2024]. In the following, we provide, as background information, a brief overview of the most salient techniques. Figure 1 provides a conceptual overview of compliance monitoring.
Figure 1: Conceptual overview of compliance monitoring with conformance checking or compliance monitoring, adapted from [klessascheckUnlocking2024, lyComplianceMonitoringBusiness2015]
In general, regulations are usually available in the form of textual documents [klessascheckReviewingConformance2024]. Hence, for monitoring regulatory process compliance, a specification, or operationalization, becomes necessary to relate the regulation to the business process under investigation. Depending on the technique and approach, which we will further elaborate upon in the following, different formal representations are created: Regulatory constraints relevant for a business process are operationalized, that is, used to create prescriptive process models or specific formal process constraints. These are subsequently used in various compliance monitoring techniques. Regardless of techniques and the type of approach (i.e., whether design-time or run-time regulatory compliance is checked with specific techniques), these operationalized regulations are intermediate, explicitly modelled, representations that encapsulate regulatory constraints, and are therefore an important component when assessing regulatory compliance of business processes [groefsemaUseConformanceCompliance2022].
2.1.1 Conformance Checking
Central to the execution of business processes are information systems, which capture information about the processes and their execution in event logs. These event logs, containing ordered records of individual process instances and the events that record what activity has been executed when [van2012process], are the basis of conformance checking: The process event data of event logs can be linked to a process model, allowing insights into the relation between observed and modelled behaviour [carmonaConformanceCheckingFoundations2022]. Commonly considered with conformance checking is the question of whether process behaviour captured in an event log conforms to prescribed (normative) behaviour. This normative behaviour is given in the form of a prescriptive process model [carmona2018conformance, groefsemaUseConformanceCompliance2022]. Process models are (mainly graphical) representations of business processes, and describe the process activities and their ordering [Weske19]. Consequently, techniques from the family of conformance checking can be used to assess the relation of recorded and prescribed process behaviour. Conformance checking techniques usually express the result of such an analysis in the form of fitness values, i.e., the degree of conformance between event log and process model, and provide lists of deviations, i.e., recorded process executions that deviate from the process models. Different techniques for conformance checking exist, mainly differing in the type of prescriptive model they use [dunzerConformanceCheckingStateoftheart2019]. Imperative techniques use imperative process models, such as BPMN [bpmnspec] or Petri nets, which describe only allowed behaviour. Declarative techniques apply declarative process models, such as Declare [pesicDECLAREFullSupport2007], which only express constraints on process behaviour, instead of explicitly specifying it.
Further, hybrid techniques use a mixture of imperative and declarative elements to prescribe behaviour of business processes [van2021conformance]. In general, prescriptive models formalize constraints towards one or more process perspectives, which are characteristics of processes relevant for conformance checking. These perspectives include the ordering of activities, temporal aspects, data and documents, and resources that execute the process [russellWorkflowPatternsDefinitive2015].
In practice, conformance checking has been identified as a suitable and useful approach for regulatory compliance monitoring [carmona2018conformance]. Here, regulations relevant for a business process are interpreted and operationalized into prescriptive process models that express compliance constraints. These prescriptive models are then used to monitor the regulatory compliance of recorded process event data in the form of event logs, with the help of conformance checking techniques. This is done either based on event logs (offline and ex-post, similar to auditing), or based on event streams (online, during run-time). According to Groefsema et al. [groefsemaUseConformanceCompliance2022], the main issue with using conformance checking for regulatory compliance monitoring lies in the fact that the prescriptive model being used must also be proven to be regulatory compliant with regulatory compliance checking. However, given that that conformance checking is indeed applied for regulatory compliance monitoring, a systematic investigation is warranted.
2.1.2 Compliance Checking
While conformance checking techniques can be used for regulatory compliance monitoring, other techniques that do not belong to the family of conformance checking exist. Here, we summarize them under the umbrella term of compliance checking, in line with [groefsemaUseConformanceCompliance2022, lyComplianceMonitoringBusiness2015]. Complex event processing (CEP) [luckhamPowerEvents2008] in particular, focussed on processing and aggregating streams of process event data, has been positioned as an enabling technique for regulatory compliance monitoring [lyComplianceMonitoringBusiness2015, knupleschAFramework2017, hashmiAreWeDone2018]: In CEP, process event data generated during process execution is collected and aggregated into higher-level events, to which monitoring rules and constraints are applied so that non-compliant behaviour can be detected [knupleschVisuallyMonitoring2015, hashmiAreWeDone2018]. In general, CEP can be used for monitoring various types of system requirements, such as user requirements or design properties regarding e.g. information flow of web services [fischerSecurityPolicyEnforcement2007]; the focus of using CEP for compliance monitoring of business processes, and thus of this study, rather lies on the application of techniques for monitoring process events that make up business process executions [lyComplianceMonitoringBusiness2015]. These business processes in turn are subject to compliance constraints; in the case of regulatory compliance monitoring, these constraints are representations of regulatory constraints. We therefore do not consider applications of approaches that focus on granularities beyond those of business process instances, such as system requirements or service requirements, e.g. SLAs, or compliance constraints that do not stem from regulations.
Regardless of the technique, in compliance checking approaches, it is also necessary to interpret compliance requirements and specify compliance constraints [lyComplianceMonitoringBusiness2015]; these constraints, similar to a prescriptive model of conformance checking techniques, relate to process perspectives and could be imperative, declarative, or hybrid representations of prescribed behaviour [lyComplianceMonitoringBusiness2015]. For example, declarative representations include the extended Compliance Rule Graph Language (eCRG) which consists of control flow, data, resources and temporal elements, with which regulatory constraints can be expressed in prescriptive models [hashmiAreWeDone2018, knupleschVisuallyMonitoring2015]; further examples include linear-temporal logic (LTL) and Declare [hashmiAreWeDone2018]. Imperative techniques for compliance checking might use, similar to conformance checking, traditional imperative process models from which then monitoring queries for use in CEP environments can be derived [weidlichEventBasedMonitoring2011]. For a more comprehensive technical overview of compliance checking, we refer to the work of Hashmi et al. [hashmiAreWeDone2018].
In the literature, compliance checking has been characterized as being more suitable for monitoring regulatory compliance than conformance checking. As argued by [hashmiAreWeDone2018, hashmiNormativeRequirementsRegulatory2016], this is because the imperative models typically used for conformance checking would primarily express business rules and not regulatory constraints [hashmiAreWeDone2018, groefsemaUseConformanceCompliance2022, hashmiNormativeRequirementsRegulatory2016]. Business rules are here understood as “declarations which constrain, derive and give conditions for existence, representing the knowledge of the business, [… and] define the conditions under which a process is carried out or the new conditions that will exist after a process has been completed” [araujoAMethod2010], whereas regulatory constraints describe conditions and obligations stemming directly from the legal domain [hashmiNormativeRequirementsRegulatory2016, hashmiAreWeDone2018]. Still, existing research also underlines similarities and synergies between conformance and compliance checking approaches [lyComplianceMonitoringBusiness2015]. Further, the difference between conformance and compliance checking approaches has been blurred due to terminological inconsistencies: in fact, checking whether observed behaviour follows a specification, is in the strictest sense conformance checking, even if the goal is to prove regulatory compliance [groefsemaUseConformanceCompliance2022]. Notably, we see conformance checking applied in practice for regulatory compliance monitoring purposes, also using declarative techniques and prescriptive models [klessascheckReviewingConformance2024]); we are further aware of studies that position themselves as checking compliance directly. Therefore, it appears prudent to further investigate similarities and differences in practical uses of compliance monitoring as reported in studies that position themselves as applying compliance and conformance checking.
2.2 Related Work
Various contributions have already aimed at systematically assessing the capabilities of existing compliance monitoring approaches, in particular compliance checking and conformance checking techniques. Ly et al. [lyComplianceMonitoringBusiness2015] present a framework for systematic comparison of compliance monitoring approaches along modelling requirements, execution requirements and user requirements covered by existing tooling. Modelling requirements include the ability to model time, data, and resource constraints, whereas user requirements include the ability to reactively or proactively detect violations, the ability to explain the root cause of deviations, and the ability to quantify the degree of compliance. Execution requirements relate to the ability of dealing with information available during execution of a business process on an event level, such as dealing with activities that are non-atomic, checking for constraints of an activity’ life cycle, or supporting multiple instances of the same constraint in a single process instance. Existing compliance monitoring approaches are classified according to whether they fulfil these requirements by applying all of them to one particular example.
Building on this framework, Rinderle-Ma et al. [rinderle-maPredictiveComplianceMonitoring2023] classify predictive compliance monitoring approaches along compliance monitoring requirements and capabilities. For doing so, they extend the requirements formulated by Ly et al. [lyComplianceMonitoringBusiness2015] with e.g. additional user requirements or data requirements specific to predictive process compliance monitoring. Similar to Ly et al. [lyComplianceMonitoringBusiness2015], they assess in how far existing literature covers the functionality requirements they elicited, and derive research directions from this assessment. Requirements towards compliance monitoring approaches are elicited from real-world case studies and literature in both studies; however, concrete applications of these approaches are not characterized and investigated for their particular properties. While the requirements are useful for describing the capabilities of compliance monitoring approaches, they are insufficient to characterize how these approaches are applied in a way that goes beyond their ability to meet e.g. modelling, execution, and user requirements. Instead, we believe more emphasis should be placed on how prescriptive models come to be, to what end uses of approaches are reported, and in what ways the results are represented (see also e.g. Rehse et al. [rehse2022process]), in particular for regulatory compliance monitoring.
Similarly, the literature on conformance checking has focussed on technical aspects exclusively, and less on conformance checking approaches. Dunzer et al. [dunzerConformanceCheckingStateoftheart2019] analyse conformance checking techniques for the modelling language and the type of algorithm being used, the kind of metric with which conformance is expressed, and the process perspective that is considered (e.g., the ordering of activities or temporal aspects). However, the work does not focus on compliance checking with regulations in particular, nor is the creation of the prescriptive models analysed. Further, Oliart et al. [OLIART2022104076] exclusively investigate approaches that assess clinical guidelines. Concretely, the guidelines that are used by existing studies are analysed per disease, their complexity is assessed, as well as the prescriptive model in terms of, e.g., nodes in the process model or number of declarative constraints — there is no further consideration of how these prescriptive models are created, nor what the results are or how the techniques are applied.
Thus, we see a need to investigate how applications of compliance monitoring use regulations in a practical sense beyond purely technical considerations, how they are applied, and what results they provide.
3 Research Method
To address the research questions above, we conduct a systematic literature review (SLR) to identify scientific studies that apply a) conformance checking techniques and b) other techniques of the compliance space, for regulatory compliance monitoring of business processes with concrete regulations. In this, our study constitutes a descriptive review, as we aim to identify trends and patterns in the application of compliance monitoring techniques [pareSynthesizingInformationSystems2015]. For this, we adhere to the eight-step method of Okoli [okoliGuideConductingStandalone2015]. These steps are as follows: (1) purpose of the literature review, (2) protocol and training, (3) searching for the literature, (4) practical screen, (5) quality appraisal, (6) data extraction, (7) synthesis of studies, and finally (8) writing the review. Additionally, we use a concept matrix as a framework for presenting the results [websterAnalyzingPrepareFuture2002].
For the purpose of this SLR, as elaborated upon in Sec. 1, we aim to assess the existing state of the art of approaches that apply regulatory compliance monitoring techniques, and to identify opportunities for further research. For the protocol and training step, we set up a shared document in which we documented all identified articles and reasons for their inclusion respective exclusion. For the search step, we conduct two separate searches using four scientific databases (ACM, IEEE Explore, Science Direct, Web of Science), up to and including May 2024. Conducting two separate searches allows us to better differentiate between studies that position themselves as applying conformance checking techniques and as applying compliance checking techniques for regulatory compliance monitoring, thus allowing us to investigate commonalities and differences between these two perspectives on compliance monitoring. Figures 2 and 3 illustrate our search process.
The first search (subsequently referred to as Search A, illustrated in Fig. 2), aims to identify contributions that apply conformance checking as a technique for regulatory compliance monitoring exclusively. The second search (referred to as Search B, shown in Fig. 3), aims to identify contributions that apply any other technique (which we subsume as compliance checking following [lyComplianceMonitoringBusiness2015, vanbeestCrossInstanceRegulatoryCompliance2023]) for regulatory compliance monitoring. Doing so allows us to differentiate between approaches that consider compliance monitoring from the conformance checking perspective, and those that come from the compliance checking perspective unrelated to conformance checking techniques.
In Search A, we explicitly focus on studies that position themselves as applying conformance checking techniques as an approach for regulatory compliance monitoring. Moreover, we include terms commonly associated with regulations. The complete search term (“conformance checking” AND (“law” OR “guideline”, “manual” OR “reference” OR “quality” OR “compliance” OR “regulation”)) is applied to the title, abstract, and keyword search, where applicable. In Web of Science, we filtered for the categories of computer science and business.
In Search B, we explicitly search for studies that position themselves as applying techniques besides conformance checking (meaning, other compliance checking techniques) for regulatory compliance monitoring. In line with Search A, we include terms commonly associated with regulations. The complete search term ((“compliance checking” OR “compliance monitoring”) AND (“law” OR “guideline” OR “manual” OR “reference” OR “quality” OR “compliance” OR “regulation”)) is applied to the title, abstract, and keyword search, where applicable. [^1] Note that for the search we use compliance monitoring and compliance checking interchangeably for any study that mentions applications of compliance monitoring or checking techniques, due to a certain flexibility of the terms’ use we observed in the literature. In Web of Science, we filtered for the categories of computer science and business.
Aiming to investigate the scientific support for practical compliance monitoring with conformance checking and other techniques, we have set up several exclusion (EX) and inclusion (IN) criteria to screen for relevant studies. In particular, we exclude (EX1) all studies that assess this topic in an abstract manner, such as literature reviews, (EX2) studies that provide general techniques or metrics for regulatory compliance monitoring without application to recorded process executions, and (EX3) studies where compliance monitoring techniques are not applied to a concrete business process and regulation. We explicitly include (IN1) studies that apply compliance monitoring techniques to business processes (meaning, recorded process executions) against specific guidelines, laws, regulations, manuals, etc., (IN2) which are peer-reviewed and written in English. [^2] We did not, however, exclude peer-reviewed studies based on the venue or journal in which they appeared, in order to capture a wide range of contributions, which is consistent with descriptive reviews [pareSynthesizingInformationSystems2015]. This constitutes the practical screen.
An initial Search A – illustrated in Fig. 2 – with the databases and search term described above produces, after deduplication, 235 candidates for analysis. Afterward, we conduct the quality appraisal by deciding, first based on the title and abstract and then on the full text, whether the found studies meet our criteria. In total, 51 remain after the title and abstract screening and 25 after the full-text screening. We add twelve papers through the authors’ expert knowledge and identify four additional studies via a “forward-backward-search” we conduct based on the relevant papers of the full-text screening to find potentially missing studies. Both steps are in line with existing recommendations for conducting literature reviews to ensure that the search covers sufficient literature [okoliGuideConductingStandalone2015, vombrockeStandingShouldersGiants2015]. This leads to 41 papers we consider relevant for our analysis from Search A.
Figure 2: Search process of Search A for relevant literature, adapted from [klessascheckReviewingConformance2024]
For Search B, illustrated in Fig. 3, we identify 851 candidates for analysis after an initial search and deduplication, as described above. Afterwards, similarly to Search A, we conduct the quality appraisal; after title and abstract screening, 119 articles remain, and 13 after full-test screening. Through a “forward-backward-search”, we identify two additional relevant studies, and two via expert knowledge. Thus, Search B has produced 17 papers we consider relevant for our analysis.
Figure 3: Search process of Search B for relevant literature, adapted from [klessascheckReviewingConformance2024]
As part of papers added via expert knowledge, we filter out studies in Search A that clearly apply compliance checking techniques instead of conformance checking techniques and add them to Search B as additional expert knowledge; similarly, studies that clearly apply conformance checking techniques instead of compliance checking are screened out and added to Search A as additional expert knowledge. This was done for five studies in total (one study from Search A to B, and four from B to A). For clarity, Figs. 2 and 3 show these studies as part of the “expert knowledge” step, and they have been already included in the preceding paragraphs. After having identified 41/17 (Search A/Search B) relevant studies, we continue with data extraction and synthesis of the extracted information for each of the two searches.
In relation to our research questions, we define a set of categories and corresponding characteristics to capture the distinctions within the compliance monitoring application of each study. We first extract detailed information for each category. Subsequently, we analyse the extracted information for similarities in their characteristics. Finally, as the synthesis step, we identify potential levels of abstraction that accurately reflect the information while avoiding excessive specificity. This approach is consistent with Okoli’s [okoliGuideConductingStandalone2015] methodology, where our synthesis step is qualitative.
Figure 4: Lifecycle and phases of compliance monitoring, adapted from Carmona et al.’s lifecycle of conformance checking projects [carmona2018conformance], used in this study to guide data extraction and synthesis
For the results of Search A and Search B, we identify characteristics in relation to the lifecycle model of conformance checking projects of Carmona et al. [carmona2018conformance], which divides the application of conformance checking into three phases: 1.) preparation, where a prescriptive model is formalized, process event data is captured in an event log, and the two related; 2.) action, where a compliance monitoring technique, such as conformance checking, is applied; and 3.) reflection, where results are interpreted and insights derived. These steps are closely related to the generalized approach of compliance monitoring described by Ly et al. [lyComplianceMonitoringBusiness2015] (where compliance rules are formally specified or modelled, applied to process event data generated by process executions, and results are reported), although more structured with an explicit iteration; therefore we adopt the conformance checking lifecycle model as our theoretical lens. Figure 4 displays the adapted lifecycle of compliance monitoring. The results, in the form of a concept matrix following Webster and Watson [websterAnalyzingPrepareFuture2002], can be found in Table 1.
Table 1: Concept matrix [websterAnalyzingPrepareFuture2002] displaying the synthesis results. PM refers to prescriptive model; the domains and results are abbreviated. [^4] The domains are Customer Services, Data Processing, Education, Finance, Healthcare, IT Service Management, Human Resources, Manufacturing, Nuclear Power, Public Services, Rent, Science, Software, Supply Chain, Trade. Besides Qualify and Quantify, the results contain Breakdown and Compare, Localize and Show, and Explain and Diagnose.
Preparation Action Reflection
Study
Domain
Search
Goal
PM Creation
PM Validated
Data Type
Data Aval.
Exp. Knowl.
Technique
Proc. Persp.
Application
Results
Reference
Domain
Technique Improv.
Process Improv.
Demonstration
manual
automated
yes
Real-world
Synthetic
Available
Not available
yes
Imperative
Declarative
Hybrid
Control Flow
Temporal
Data
Resources
Manual
Tool-supported
Automated
Real-Time
Qualify
Quantify
BD & Comp.
Loc. & Show
Expl. & Diagn.
[accorsiExploitationProcessMining2012] Fin A
[baumgrassConformanceCheckingRBAC2012] Fin A
[bottrighiConformanceCheckingExecuted2012] HC A
[buijsCrossOrganizationalProcessMining2012] PS A
[chesaniAbducingComplianceIncomplete2016] PS A
[chikkaConformanceCheckingMethodology2020] HC A
[GerkeTamm2014] ITSM A
[gerkeMeasuringComplianceProcesses2009] ITSM A
[huangOnlineTreatmentCompliance2014] HC A
[jansBusinessProcessMining2011] Fin A
[jansFieldStudyUse2014] Fin A
[lemosUsingProcessMining2011] SW A
[lenkowiczAssessingConformityClinical2018] HC A
[liangIdentificationHumanOperation2021] Nuc A
[piovesanTemporalConformanceAnalysis2018] HC A
[placidiProcessMiningOptimize2021] HC A
[premchaiswadiProcessModelingBottleneck2015] Sci A
[rinnerProcessMiningConformance2018] HC A
[rovaniDeclarativeProcessMining2015] HC A
[sangilHeuristicsBasedProcessMining2020] PS A
[sarnoHybridAssociationRule2015] Fin A
[satoConformanceCheckingDifferent2020] HC A
[spiottaTemporalConformanceAnalysis2017] HC A
[stertzRoleTimeData2021] Man A
[taghiabadiComplianceCheckingDataAware2014] HC,Fin A
[xuModelingClinicalActivities2020] HC A
[xuApplyingClinicalGuidelines2020] HC A
[zamanInnovativeOnlineProcess2019] DP A
[zamanEnablingGDPRCompliance2020] DP A
[beneventoProcessModeling2023] HC A
[grugerVerifyingGuideline2022] HC A
[grugerWeightedViolations2023] HC A
[savinoProcessMining2023] HC A
[grugerDeclarativeGuideline2023] HC A
[ramezani2012WhereDidI] HC A
[breitmayerPermissionAnalysisObjectCentric2024] Edu A
[savinoErrorCorrectingMethodologyEvaluating2024] HC A
[zhangReorderedFuzzyConformance2024] HC A
[salnitriSecuritybyDesignIdentificationSecurityCritical2018] Fin A
[wangRegulatorySupervisionComputational2018a] Sup A
[wangAnalyzingTransactionCodes2018a] Man A
[knupleschAFramework2017] HR B
[lyComplianceMonitoringBusiness2015] HC B
[vanbeestCrossInstanceRegulatoryCompliance2023] Sup B
[araujoAMethod2010] Rent B
[jiang2014RegulatoryCompliance] Trade B
[jiangComplianceChecking2015] Trade B
[weidlichEventBasedMonitoring2011] CS B
[hashmiNormativeRequirementsBusiness2014] PS B
[WOS:000540648600033] DP B
[bukhshComplianceCheckingShipment2017] Trade B
[caronBusinessRulePatterns2013] Fin,HC B
[gongBpCMonRulebasedMonitoring2017] HC B
[gonzalezComplianceRequirements2021] PS B
[hoangChallengesDiscoveringPatient2022] HC B
[lyMonitoringBusinessProcess2011] Fin B
[middletonFrameworkContinuousCompliance2009] HC B
[WOS:000424457700006] Fin B
For the preparation phase, we capture the process domain, the goal with which a compliance monitoring technique is applied (to improve compliance monitoring techniques, to improve the business process compliance, or to demonstrate compliance monitoring techniques), as well as whether the prescriptive model is created manually or automatically, and whether it is explicitly validated through e.g., process experts or stakeholders. We also capture whether the process execution data in the form of event logs has synthetic or real-world origins, and whether it is publicly available or not.
For the action phase, we identify whether expert knowledge is necessary for applying compliance monitoring techniques, whether imperative, declarative, or hybrid techniques [^5] This we determine primarily based on the type of prescriptive model — i.e., declarative, imperative, or hybrid — since the exact technique for regulatory compliance monitoring was not made transparent in all studies we investigated. Nonetheless, the way in which compliance constraints are formalized appears to be a good indicator of the type of technique applied to them. are used (see [carmona2018conformance, de2015alignment]), which process perspectives are assessed (control flow, temporal, data, or resource, see [russellWorkflowPatternsDefinitive2015]), and whether the compliance monitoring techniques are applied manually, in a tool-supported manner, automatically, or in real-time (i.e., online). In contrast to [lyComplianceMonitoringBusiness2015], we also explicitly include the control flow process perspective. However, we do so on a higher level of abstraction than the control flow requirements formulated by [rinderle-maPredictiveComplianceMonitoring2023], which differentiate control flow constraints between existence, absence, and ordering constraints. Subsuming these under the control flow perspective, as in [dunzerConformanceCheckingStateoftheart2019], has been done to describe the application of regulatory compliance monitoring techniques more broadly.
For the reflection phase, we capture the different types of results and their representation (i.e., whether compliance is qualified, quantified, broken down and compared, or explained and diagnosed, following a classification of representations identified from current process mining tools [rehse2022process]). Here, we focus on understanding why and how results of regulatory compliance monitoring are presented, as opposed to determining whether existing approaches are able to provide specific insights to users [lyComplianceMonitoringBusiness2015, rinderle-maPredictiveComplianceMonitoring2023]. Doing so, we find that qualifying the overall degree of regulatory compliance is also relevant in concrete applications, something not explicitly covered by related work [lyComplianceMonitoringBusiness2015, rinderle-maPredictiveComplianceMonitoring2023].
For Search A, the screening, data extraction, and synthesis steps were performed by both authors and a research assistant. The initial papers found in Search A were distributed among them for further analysis as described above, and the results of that analysis were combined in a discussion. During this discussion, unclarities and disagreements were resolved. For Search B, the main author and another research assistant performed the screening, data extraction, and synthesis steps in a similar manner, while re-using the coding scheme developed during Search A. Again, they combined the results of their analyses in a discussion, and resolved unclarities and disagreements. Finally, the main author performed an assessment of the entire result of Search A and B, and discussed final changes to the analysis among the author team, until a consensus was reached. In Table LABEL:tab:categories_explained, we provide a detailed description of the categories and characteristics we extracted in Table 1. In the following, we analyse the resulting synthesis along its dimensions, thereby writing the review.
4 Results
After inductive coding and synthesizing the 58 (41 of Search A, 17 of Search B) relevant studies, we analyse the resulting data, in order to characterize their compliance monitoring mechanisms and their usages. This is done along the three phases of conformance checking projects [carmona2018conformance], being preparation (Sec. 4.1), action (Sec. 4.2), and reflection (Sec, LABEL:sec:results-reflection). We also describe commonalities and differences between the uses found in the two searches. Generally, we can observe, as shown in Fig. 5, that we find far fewer studies that position themselves as compliance checking than those framed as conformance checking studies. We also note that while conformance checking studies were published quite early, they appear to peak slightly later than compliance checking studies, which might hint at a shift in research interest or a change in the positioning of the studies; similarly, compliance checking studies seem to tail off after having reached a peak in 2017.
2009200920102010201120112012201220132013201420142015201520162016201720172018201820192019202020202021202120222022202320232024202402244662202255033331111661166331144331111220112222033111101111110YearStudiesConformanceCompliance
Figure 5: Distribution of studies found for Search A (conformance) and Search B (compliance) per year
4.1 Preparation
Domain
First, we analyse the domains in which conformance checking and compliance checking techniques are applied for regulatory compliance monitoring. Figure 6 shows an overview of the domains we identified.
CSDPEduFinHCHRITSMManNucPSRentSciSupSWTrade0101020200221177212102222113301111110111103355110002211011033DomainStudiesConformanceCompliance
Figure 6: Distribution of domains found in the studies identified with Search A (conformance) and Search B (compliance)
The majority of conformance checking studies focusses on the area of healthcare (21 studies), with finance being second (seven studies). Other areas, such as public services (three studies) or data processing (two studies), are also present. The prevalence of healthcare and finance seems to indicate that regulations in these areas are well-suited to be assessed with conformance checking.
The set of compliance checking studies, on the other hand, shows a broader distribution of domains (including one study in a healthcare setting): From healthcare, human resources, rent, public services to trade, supply chains and customer service, seven different domains are covered. This seems to suggest a broad applicability of compliance checking in many areas, although we observe a slight focus on healthcare-related studies (five times), trade and supply chain (four times), as well as finance and public services (three times, two times).
Goals
Second, we consider the goals of the studies we analysed. An overview of our findings is provided in Fig. 7.
TechniqueImprovementProcessImprovementDemonstration010102020303028281616212110102277GoalStudiesConformanceCompliance
Figure 7: Distribution of goals found in the studies identified with Search A (conformance) and Search B (compliance)
Regarding the explicitly stated goals of the studies for applying conformance checking techniques, we see that those in the area of healthcare mainly aim to improve the conformance checking technique and its application (15 of 21) or to demonstrate the applicability of conformance checking (eleven of 21), whereas only around a third of the studies making use of healthcare-related regulations aim to improve the process itself. Similarly, no finance-related contribution aims to provide actual process improvement, but all of them aim to improve conformance checking techniques, and two out of seven explicitly intend to demonstrate them. This tendency is noticeably present across all domains we identified, with more than two thirds (28) aiming to improve techniques, more than half aiming to demonstrate them (21), and less than half of the studies (16) aiming to investigate actual process improvement.
For compliance checking applications, the situation is slightly: more than half (i.e., ten of 17) of the studies aim to improve compliance checking techniques and their application, only two of 17 studies aim to improve the business process itself, and little over two fifths (i.e., seven of 17 studies) aim to demonstrate the application of compliance checking. Notably, all three studies that explicitly set out to improve process compliance are situated in the area of trade and supply chains.
Prescriptive Model Creation
Third, we consider how prescriptive models are created in the relevant studies. In all conformance checking and compliance checking studies, we see that no contribution out of the 58 identified studies automatically creates them based on the relevant regulations. In fact, all studies rely on manual input for creation of the required prescriptive model. Notably, the exact operationalization that led to the model is largely not described in detail, and instead, only descriptions or visualizations of them are provided. For example, in [accorsiExploitationProcessMining2012], a BPMN model is created manually with and validated by process experts (in this case, bank managers), and further requirements are provided informally by bank managers and auditors, and manually translated into LTL formulae by the authors.
Prescriptive Model Validation
Additionally, we note that in less than one third of the conformance checking studies (13 of 41), the prescriptive model is explicitly validated by the authors in conjunction with stakeholders or process experts. [^6] The remaining studies did not mention whether the prescriptive model was evaluated or not. Hence, we only capture whether an explicit evaluation took place. Interestingly, around half (seven of 13) of those studies in which a validation is reported are in the healthcare domain. We assume that this may be because of the need for expert involvement in healthcare settings to even operationalize the treatment guidelines, so that validation is a comparatively trivial addition.
In contrast to this, only two of the 17 compliance checking studies explicitly mention a validation of the prescriptive model that has been used. [^7] As above, we only noted explicit evaluations, since the remaining seven studies did not mention whether an evaluation took place or not. This may be linked to the prevalent technique improvement goal we see in compliance checking studies: in order to improve compliance checking techniques, the prescriptive models may not need to be evaluated since an evaluated prescriptive model is less relevant here than for demonstrating techniques or improving actual process compliance.
Data Origin and Data Availability
For the final aspect of the preparation phase, we look at the origin and availability of the process event data used in the analysed studies. In Fig. 8, we provide an overview of our findings.
Real-worldSynthetic1010202030304040353566121266OriginStudiesConformanceCompliance
(a) Data Origin
AvailableNot Available1010202030304040883434441313OriginStudies
(b) Data Availability
Figure 8: Data Origin and data availability for the studies identified with Search A (conformance) and Search B (compliance)
As to the conformance checking studies, 35 of 41 conformance checking studies use real-world data, although only in seven cases this data is available. Usually, available real-world process event data stems from the Business Process Intelligence Challenge (BPIC) (e.g., [taghiabadiComplianceCheckingDataAware2014]), but also from data provided by a governmental data sharing portal [sangilHeuristicsBasedProcessMining2020] or a publicly available clinical database [chikkaConformanceCheckingMethodology2020]. 21 of 35 applications with real-world data are in the healthcare domain, which is also the most represented domain in the conformance checking studies. On the other hand, only six of 41 studies use purely synthetic data, which is made available in one case. We therefore observe a low overall availability of process event data in the conformance checking space but a high degree of real-world data.
In the compliance checking studies, we see real-world process event data being used in twelve of 17 studies, and synthetic data in six of 17 studies. However, only in four of the real-world studies, this data is actually available – notably, three of them employ BPIC event logs which have been made available for research purposes beforehand [lyComplianceMonitoringBusiness2015, vanbeestCrossInstanceRegulatoryCompliance2023, gongBpCMonRulebasedMonitoring2017].
4.2 Action
Technique
The conformance techniques used in the investigated contributions, are largely imperative, with less than a third (i.e., eleven of 41) declarative and only two hybrid techniques. The largest share of declarative techniques is present in the healthcare domain (eight of eleven), underlining that the nature of clinical guidelines may be especially suited for declarative techniques. For example, [bottrighiConformanceCheckingExecuted2012] translates a clinical guideline into a formal representation based on logical programming that combines declarative and imperative aspects. Based on a Prolog implementation of the Event Calculus, they can assess the regulatory compliance of recorded process executions w.r.t. a clinical guideline. As another example, [sangilHeuristicsBasedProcessMining2020] uses ProM to compare the activities and timings of public procurement data provided by a government’s open data portal with the activity orderings and timings prescribed by relevant regulation.
For compliance checking techniques, we observe almost exclusively declarative applications (15 out of 17), regardless of domain. For example, [bukhshComplianceCheckingShipment2017] use ProM’s LTL checker to assess the regulatory compliance of a company’s trade and customs procedures, for which they manually create formal regulatory constraints from relevant statements of the regulations. The prescriptive model is imperative in nature only in two studies, for example in [weidlichEventBasedMonitoring2011] where it serves to derive event queries for use with CEP. The overall prevalence of declarative applications may be linked to mechanisms used for regulatory compliance monitoring, such as CEP, where instead of alignments with an imperative process model, done so in imperative conformance checking techniques [carmona2018conformance], standing queries are evaluated against observed process event data [weidlichEventBasedMonitoring2011] to determine compliance or non-compliance.
Process Perspective
As to the process perspective, we provide an overview of our findings in Fig. LABEL:fig:process_perspective.